Gucchi:

Which statement is true about the amount of refraction of light waves as they pass through a material, or medium? Depends on the temperature of the light wave Decreases with an increase in density of medium Increases with an increase in density of medium Depends on the angle at which the wave strikes the medium

Gucchi:

Am I right with D on this one, i'm confused

Gucchi:

@AZ

AZ:

Refraction formula \( n_1 \sin \theta_1 = n_2 \sin \theta_2\) where \( n_1\) is the incident index \( n_2\) is the refracted index \(\theta_1\) is the incident angle \(\theta_2\) is the refracted angle I don't know if you've learned the formula yet but yes you're correct

Gucchi:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ Refraction formula \( n_1 \sin \theta_1 = n_2 \sin \theta_2\) where \( n_1\) is the incident index \( n_2\) is the refracted index \(\theta_1\) is the incident angle \(\theta_2\) is the refracted angle I don't know if you've learned the formula yet but yes you're correct \(\color{#0cbb34}{\text{End of Quote}}\) Okay, thanks for this

Gucchi:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @kittybasil Yep, you're correct. Here's an explanation (also below): \(\color{#0cbb34}{\text{End of Quote}}\) Thank you

AZ:

You're welcome!

AZ:

A user messaged me and informed me that D was incorrect and upon further research I would have to revise my answer So the formula does have angle measures in it, however that's not what determines the degree to which it refracts. Refraction is the bending of light that occurs when you move from one medium to another medium. It happens because of the change in densities

Gucchi:

Hmmm, okay

Gucchi:

@AZ so would it be C?

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Gucchi @AZ so would it be C? \(\color{#0cbb34}{\text{End of Quote}}\) Correct, it would be C

AZ:

Yeah Light travels more slowly in a denser medium

Gucchi:

Okay thanks for helping me correct the answer

AZ:

No problem! Credit goes to @carmelle for catching the mistake :)

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ No problem! Credit goes to @carmelle for catching the mistake :) \(\color{#0cbb34}{\text{End of Quote}}\) yw :)